# Engineering Spatial and Molecular Features from Cellular Niches to Inform Predictions of Inflammatory Bowel Disease

### v1

@@ Abstract
Differentiating between the two main subtypes of Inflammatory Bowel Disease (IBD): Crohn’s disease (CD) and ulcerative colitis (UC) is a persistent clinical challenge due to overlapping presentations. This study introduces a novel computational framework that employs spatial transcriptomics (ST) to create an explainable machine learning model for IBD classification. We analyzed ST data from the colonic mucosa of healthy controls (HC), UC, and CD patients. Using Non-negative Matrix Factorization (NMF), we first identified four recurring cellular niches, representing distinct functional microenvironments within the tissue. From these niches, we systematically engineered 44 features capturing three key aspects of tissue pathology: niche composition, neighborhood enrichment, and niche-gene signals. A multilayer perceptron (MLP) classifier trained on these features achieved an accuracy of $0.774\pm 0.161$ for the more challenging three-class problem (HC, UC, and CD) and $0.916\pm 0.118$ in the two-class problem of distinguishing IBD from healthy tissue. Crucially, model explainability analysis revealed that disruptions in the spatial organization of niches were the strongest predictors of general inflammation, while the classification between UC and CD relied on specific niche-gene expression signatures. This work provides a robust, proof-of-concept pipeline that transforms descriptive spatial data into an accurate and explainable predictive tool, offering not only a potential new diagnostic paradigm but also deeper insights into the distinct biological mechanisms that drive IBD subtypes.
Keywords:
cellular niches explainable machine learning (XML) feature engineering inflammatory bowel disease spatial transcriptomics.
@@ 1 Introduction
Inflammatory Bowel Disease (IBD) presents a significant clinical challenge due to the diagnostic and therapeutic ambiguity between its two main subtypes, Crohn’s disease (CD) and ulcerative colitis (UC) [1, 2]. Though pathologically distinct, overlapping clinical presentations can complicate diagnosis, which is critical as treatment strategies diverge substantially [3]. Patient heterogeneity in presentation and therapeutic response underscores the need for precise diagnostic tools reflecting underlying cellular and molecular complexity [4].
Single-cell RNA sequencing (scRNA-seq) has provided unprecedented insight into this complexity, revealing diverse immune, stromal, and epithelial cell states in the inflamed gut [5, 6, 7]. Garrido-Trigo et al. combined scRNA-seq with spatial imaging to highlight that the greatest inter-patient variability in IBD lies within myeloid cells, particularly in macrophages and neutrophils [8]. While essential for unraveling disease mechanisms at a cell-by-cell resolution [9], dissociation-based scRNA-seq loses native tissue architecture. Spatial transcriptomics (ST) preserves this, enabling gene expression analysis in its morphological context [10, 11]. Here, we use ST to investigate whether the organization of cells into functional microenvironments can serve as a robust diagnostic tool for IBD subtypes.
We hypothesized that disruptions in the structure of cellular niches, which are localized communities of interacting cells, are distinct hallmarks of UC and CD. Our computational approach classified cell types [12] to map cell populations across ST data from colonic tissue, then applied non-negative matrix factorization (NMF) to identify four recurring cellular niches defined by unique cell types combinations. This analytical framework, which deconstructs spatial data into functional units, has proven effective in identifying pathological microenvironments in other complex inflammatory conditions like idiopathic pulmonary fibrosis [13]. Similar ST approaches have mapped healing programs in mouse models of colitis [14], fibrosis-associated networks in stricturing CD [15], and cellular ecosystems correlating with UC therapeutic response [16]. To build a predictive model, we engineered 44 features capturing three key aspects of the tissue: niche composition (relative abundance), neighborhood enrichment (spatial interactions), and niche-gene signals (localized gene expression). A multilayer perceptron (MLP) classifier was then trained on these features to distinguish between healthy controls (HC), UC, and CD.
@@ 2.1 Data
The study used two publicly available datasets from the NCBI Gene Expression Omnibus (GEO). The primary dataset, GSE234713 [8, 17], consisted of ST data from colonic mucosa. This data was collected from nine formalin-fixed paraffin-embedded (FFPE) human samples, including three healthy non-IBD controls (HC), three patients with CD, and three with UC. Generated using the NanoString CosMx [18] platform, this dataset provided expression profiles for 980 genes. The secondary dataset, GSE214695 [8, 19], contained scRNA-seq data from a separate cohort of healthy and IBD colonic mucosa samples. These scRNA-seq data served as a reference for classifying and assigning cell types within the primary spatial dataset. A detailed breakdown of the dataset, including the conditions (health or disease states), FFPE samples per state, number of fields of view (FOV) per sample, and number of cells per sample, is provided in Table 1.
@@ 2.4.2 Niche neighborhood enrichment score
To capture local spatial organization, we examined cell–cell neighborhoods within each FOV. A KD-tree [22] was used to identify neighbors, where any cell within a distance equal to twice the diameter of the focal cell was considered a neighbor (Figure 1). Let $O_{i,j}$ be the observed count of ordered neighbor pairs in which the focal cell belongs to Niche $i$ and the neighboring cell to Niche $j$ ($i,j\in\{1,2,3,4\}$). Let $p_{i}$ represent the overall proportion of Niche $i$ cells in the FOV and let $T=\sum_{i=1}^{4}\sum_{j=1}^{4}O_{i,j}$ be the total number of ordered pairs. The expected count under random mixing is $E_{i,j}=T\cdot p_{i}p_{j}$. A Laplace-smoothed log-ratio was then calculated to quantify enrichment:
@@ 2.5 Multilayer perceptron
A MLP [27, 28] was trained to classify each field of view (FOV) into one of three condition groups: HC, UC, or CD. The model input consisted of 44 engineered features ($p=44$), and the output was a three-class categorical variable ($k=3$). A pipeline was constructed to standardize features followed by MLP classification. Hyperparameter optimization [29] was performed using randomized search over 30,000 candidate configurations with a three-fold stratified group cross-validation scheme ($n=171$ FOVs). The search explored activation functions ($\mathrm{ReLU}$ [30] or $\tanh$ [31]), regularization strengths $\alpha\sim\mathrm{LogUniform}(10^{-5},10^{-1})$, batch sizes $\{2,4,8,16\}$, and seven hidden layer architectures: $(25)$ [single small layer], $(32,16,8)$ [progressively tapering], $(40,20,10,5)$ [deeper with small neurons], $(44,22)$ [starting at feature count, then tapering], $(50,25,12)$ [moderately wide tapering], $(50,50)$ [uniform width], and $(64,32)$ [wide, shallow network].
The optimal architecture (Figure 2) consisted of four hidden layers with the number of neurons decreasing in size: $40\rightarrow 20\rightarrow 10\rightarrow 5$. It used the ReLU activation function, a batch size of 4, and an $L^{2}$ regularization [32] parameter $\alpha=0.0010912668217800472$. This configuration achieved a mean F1-score of 0.712 for the three-class problem.The model was optimized using the Adam solver [33] with an adaptive learning rate and a maximum of 1000 iterations. Performance was evaluated using weighted F1-score as the primary metric, with additional reporting of accuracy, precision, and recall [34]. Final evaluation was based on mean performance across folds and included a confusion matrix to assess misclassification patterns across condition groups.
@@ 2.6.1 Causal discovery
For each dataset, FCI-Stable was run with a significance level of $\alpha=0.05$ and the orientation rule set to maxp [39]. This algorithm identifies a partially directed acyclic graph (PDAG), where nodes represent either spatial niches, enrichment variables, or niche–gene features, and edges encode potential direct or conditional dependencies. Analyses were performed separately for each condition group and for the combined disease variable, yielding graphical models that capture directional relationships between features and disease. The resulting causal graphs were exported as SIF files for downstream visualization and interpretation.
@@ 2.6.2 Statistical tests
Neighborhood enrichment was quantified for each FOV by comparing observed versus expected counts of adjacent niche interactions. These enrichment scores were then analyzed to determine whether the spatial organization of niches differed across HC, UC, and CD. A global Kruskal–Wallis test [40] was first performed for each niche interaction to evaluate overall differences among the three disease groups. When a global test reached significance ($p<0.05$), pairwise Mann–Whitney U tests [24, 25] were subsequently conducted to identify specific group-level differences. The Benjamini–Hochberg procedure [26] was again used to correct for multiple comparisons. Significant pairwise comparisons were reported with their adjusted $p$-values, and the relative direction of change (“Up” or “Down”) indicated whether the enrichment of a given niche interaction was higher or lower in one group compared to another.
@@ 2.6.3 Feature importance analysis
To interpret the trained MLP classifier, permutation importance (PI) [41, 42] was computed to quantify the contribution of each feature to model performance. For each feature, its values were randomly permuted across samples while keeping other features constant, and the resulting decrease in weighted F1-score was recorded. This procedure was repeated multiple times to estimate the mean and standard deviation (SD) of the F1-score change.
@@ 3.2.2 Niche–gene features with information-theoretic selection
A total of 24 niche-gene features were identified (Table 2). HC-associated features were related to the genes PIGR, IGHG2, HLA-DRB1, IGHG1, CD38, STAT1, and CHI3L1 expressed in Niche 1, 2, and 4 cells; while UC-associated features were related to the genes IGKC, IGFBP5, HLA-DQB1, CASP3, and COL3A1 expressed in Niche 2, 3, and 4 cells; and CD-associated features were related to the genes MZT2A, COL9A2, FZD1, HDAC5, COL5A1, SOX6, FOXF1 expressed in Niche 3 and 4 cells. These features were integrated with niche composition and neighborhood enrichment features to form the final 44-feature dataset used for classification.
@@
Niche-Gene Feature MI Score Condition Group Adj. p-value
niche_1_gene_HLA-DRB1 0.351 HC 0.037
niche_1_gene_IGHG2 0.351 HC <0.001
niche_1_gene_PIGR 0.357 HC <0.001
niche_2_gene_IGFBP5 0.271 UC <0.001
niche_2_gene_IGHG1 0.348 HC <0.001
niche_2_gene_IGHG2 0.352 HC <0.001
niche_3_gene_COL9A2 0.230 CD 0.020
niche_3_gene_HLA-DQB1 0.261 UC <0.001
niche_3_gene_IGFBP5 0.276 UC <0.001
niche_3_gene_MZT2A 0.244 CD <0.001
niche_4_gene_CASP3 0.280 UC <0.001
niche_4_gene_CD38 0.401 HC <0.001
niche_4_gene_CHI3L1 0.363 HC <0.001
niche_4_gene_COL3A1 0.267 UC <0.001
niche_4_gene_COL5A1 0.252 CD <0.001
niche_4_gene_FOXF1 0.211 CD 0.011
niche_4_gene_FZD1 0.352 CD <0.001
niche_4_gene_HDAC5 0.262 CD 0.004
niche_4_gene_HLA-DQB1 0.277 UC <0.001
niche_4_gene_IGHG2 0.369 HC <0.001
niche_4_gene_IGKC 0.313 UC 0.001
niche_4_gene_MZT2A 0.209 CD 0.002
niche_4_gene_SOX6 0.235 CD 0.010
niche_4_gene_STAT1 0.367 HC <0.001
@@ 3.3 Multilayer perceptron
Figure 5 shows the confusion matrices for both classification tasks. In the three-class case (Figure 5(a)), HC samples were classified perfectly, while misclassifications primarily occurred between UC and CD, indicating overlapping features between these two disease states. In the two-class case (Figure 5(b)), the model showed strong separation between HC and IBD, with only a small number of misclassifications.
The classification reports are provided in Tables 3, 4, and 5. For the more challenging three-class problem, the macro-averaged F1-score was 0.741 (accuracy $0.774\pm 0.161$), with perfect HC classification and errors between UC and CD, reflecting their biological similarity. In comparison, the two-class problem showed excellent separation of HC and IBD, with overall accuracy of $0.916\pm 0.118$.
Overall, these results indicate that while distinguishing between UC and CD remains challenging due to biological similarity, the MLP classifier is highly effective at differentiating IBD patients from HC.
@@ 3.4.1 Causal discovery
Figure 6: Causal graphs depicting relationships between features and condition (Disease/Health State): (a) Niche composition, (b) Enrichment score, and (c) Niche-gene features.
@@ 3.4.2 Statistical tests
Table 6: Pairwise comparisons of niche compositions across disease groups.
NMF Factor Group 1 Group 2
Difference
Between
Means
Direction* p-value Observation
@@
Table 7: Pairwise comparisons of niche interactions across disease groups.

### v2

@@ Abstract
Differentiating between the two main subtypes of Inflammatory Bowel Disease (IBD): Crohn’s disease (CD) and ulcerative colitis (UC) is a persistent clinical challenge due to overlapping presentations. This study introduces a novel computational framework that employs spatial transcriptomics (ST) to create an explainable machine learning model for IBD classification. We analyzed ST data from the colonic mucosa of healthy controls (HC), UC, and CD subjects. Using Non-negative Matrix Factorization (NMF), we first identified four recurring cellular niches, representing distinct functional microenvironments within the tissue. From these niches, we systematically engineered 44 features capturing three key aspects of tissue pathology: niche composition, neighborhood enrichment, and niche-gene signals. A multilayer perceptron (MLP) classifier trained on these features achieved an accuracy of $0.774\pm 0.161$ for the more challenging three-class problem (HC, UC, and CD) and $0.916\pm 0.118$ in the two-class problem of distinguishing IBD from healthy tissue. Crucially, model explainability analysis revealed that disruptions in the spatial organization of niches were the strongest predictors of general inflammation, while the classification between UC and CD relied on specific niche-gene expression signatures. This work provides a robust, proof-of-concept pipeline that transforms descriptive spatial data into an accurate and explainable predictive tool, offering not only a potential new diagnostic paradigm but also deeper insights into the distinct biological mechanisms that drive IBD subtypes.
@@ 1 Introduction
Inflammatory Bowel Disease (IBD) presents a significant clinical challenge due to the diagnostic and therapeutic ambiguity between its two main subtypes, Crohn’s disease (CD) and ulcerative colitis (UC) [1, 2]. Though pathologically distinct, overlapping clinical presentations can complicate diagnosis, which is critical as treatment strategies diverge substantially [3]. Subject heterogeneity in presentation and therapeutic response underscores the need for precise diagnostic tools reflecting underlying cellular and molecular complexity [4].
Single-cell RNA sequencing (scRNA-seq) has provided unprecedented insight into this complexity, revealing diverse immune, stromal, and epithelial cell states in the inflamed gut [5, 6, 7]. Garrido-Trigo et al. combined scRNA-seq with spatial imaging to highlight that the greatest inter-patient variability in IBD lies within myeloid cells, particularly macrophages and neutrophils [8]. While essential for unraveling disease mechanisms at a cell-by-cell resolution [9], dissociation-based scRNA-seq loses native tissue architecture. Spatial transcriptomics (ST) preserves this, enabling gene expression analysis in its morphological context [10, 11]. Here, we use ST to investigate whether the organization of cells into functional microenvironments can serve as a robust diagnostic tool for IBD subtypes.
We hypothesized that disruptions in the structure of cellular niches [^1] In this article, the term niche is used to mean a latent factor. This is a data-derived unit representing recurring patterns of cellular organization inferred from NMF, rather than a predefined biological microenvironment., which are localized communities of interacting cells, are distinct hallmarks of UC and CD. Our computational approach classified cell types [12] to map cell populations across ST data from colonic tissue, then applied non-negative matrix factorization (NMF) to identify four recurring cellular niches defined by unique cell type combinations. This analytical framework, which deconstructs spatial data into functional units, has proven effective in identifying pathological microenvironments in other complex inflammatory conditions like idiopathic pulmonary fibrosis [13]. Similar ST approaches have mapped healing programs in mouse models of colitis [14], fibrosis-associated networks in stricturing CD [15], and cellular ecosystems correlating with UC therapeutic response [16]. To build a predictive model, we engineered 44 features capturing three key aspects of the tissue: niche composition (relative abundance), neighborhood enrichment (spatial interactions), and niche-gene signals (localized gene expression). A multilayer perceptron (MLP) classifier was then trained on these features to distinguish between healthy controls (HC), UC, and CD.
@@
Given the methodological complexity of this article, we provide an overview of the computational pipeline in Figure 7 (Appendix).
@@ 2.1 Data
The study used two publicly available datasets from the NCBI Gene Expression Omnibus (GEO). The primary dataset, GSE234713 [8, 17], consisted of ST data from colonic mucosa. This data was collected from nine formalin-fixed paraffin-embedded (FFPE) human samples, including three healthy non-IBD controls (HC), three subjects with CD, and three with UC. Generated using the NanoString CosMx [18] platform, this dataset provided expression profiles for 980 genes. The secondary dataset, GSE214695 [8, 19], contained scRNA-seq data from a separate cohort of healthy and IBD colonic mucosa samples. These scRNA-seq data served as a reference for classifying and assigning cell types within the primary spatial dataset. A detailed breakdown of the dataset, including the conditions (health or disease states), FFPE samples per state, number of fields of view (FOV) per sample, and number of cells per sample, is provided in Table 1.
@@ 2.4.2 Niche neighborhood enrichment score
To capture local spatial organization, we examined cell–cell neighborhoods within each FOV. A KD-tree [22] was used to identify neighbors, where any cell within a distance equal to twice the diameter of the focal cell was considered a neighbor (Figure 1). Let $O_{i,j}$ be the observed count of ordered neighbor pairs in which the focal cell belongs to Niche $i$ and the neighboring cell to Niche $j$ ($i,j\in\{1,2,3,4\}$). Let $p_{i}$ represent the overall proportion of Niche $i$ cells in the FOV and let $T=\sum_{i=1}^{4}\sum_{j=1}^{4}O_{i,j}$ be the total number of ordered pairs. The expected count under random mixing is $E_{i,j}=T\cdot p_{i}p_{j}$. A Laplace-smoothed log-ratio was then calculated to quantify enrichment $\mathrm{S}_{i,j}$:
@@ 2.5 Multilayer perceptron
A MLP [27, 28] was trained to classify each field of view (FOV) into one of three condition groups: HC, UC, or CD. The model input consisted of 44 engineered features ($p=44$), and the output was a three-class categorical variable ($k=3$). A pipeline was constructed to standardize features followed by MLP classification. Hyperparameter optimization [29] was performed using randomized search over 30,000 candidate configurations with a three-fold stratified group cross-validation scheme ($n=171$ FOVs) with subject ID as the grouping variable, i.e., all FOVs from a given subject were assigned to the same fold. The search explored activation functions ($\mathrm{ReLU}$ [30] or $\tanh$ [31]), regularization strengths $\alpha\sim\mathrm{LogUniform}(10^{-5},10^{-1})$, batch sizes $\{2,4,8,16\}$, and seven hidden layer architectures: $(25)$ [single small layer], $(32,16,8)$ [progressively tapering], $(40,20,10,5)$ [deeper with small neurons], $(44,22)$ [starting at feature count, then tapering], $(50,25,12)$ [moderately wide tapering], $(50,50)$ [uniform width], and $(64,32)$ [wide, shallow network].
The optimal architecture (Figure 2) consisted of four hidden layers with the number of neurons decreasing in size: $40\rightarrow 20\rightarrow 10\rightarrow 5$. It used the ReLU activation function, a batch size of 4, and an $L^{2}$ regularization [32] parameter $\alpha=0.001$ (rounded; results not sensitive to finer precision). This configuration achieved a mean F1-score of 0.712 for the three-class problem. The model was optimized with Adam [33] using MLPClassifier in scikit-learn with learning_rate=’adaptive’, which reduces the base learning rate when training plateaus. This is the scheduler implemented by MLPClassifier and is distinct from the per-parameter adaptation inherent to Adam [33]. A maximum of 1000 iterations was used. Performance was evaluated using weighted F1-score as the primary metric, with additional reporting of accuracy, precision, and recall [34]. Final evaluation was based on mean performance across folds and included a confusion matrix to assess misclassification patterns across condition groups.
@@ 2.6.1 Causal discovery
For each dataset, FCI-Stable was run with a significance level of $\alpha=0.05$ and the orientation rule set to maxp [39]. This algorithm identifies a partial ancestral graph (PAG), where nodes represent either spatial niches, enrichment variables, or niche–gene features, and edges encode potential direct or conditional dependencies. Analyses were performed separately for each condition group and for the combined disease variable, yielding graphical models that capture directional relationships between features and disease. The resulting causal graphs were exported as SIF files for downstream visualization and interpretation.
@@ 2.6.2 Statistical tests
Neighborhood enrichment was quantified for each FOV by comparing observed versus expected counts of adjacent niche interactions. These enrichment scores were then analyzed to determine whether the spatial organization of niches differed across HC, UC, and CD. A global Kruskal–Wallis test [40] was first performed for each niche interaction to evaluate overall differences among the three condition groups. When a global test reached significance ($p<0.05$), pairwise Mann–Whitney U tests [24, 25] were subsequently conducted to identify specific group-level differences. The Benjamini–Hochberg procedure [26] was again used to correct for multiple comparisons. Significant pairwise comparisons were reported with their adjusted $p$-values, and the relative direction of change (“Up” or “Down”) indicated whether the enrichment of a given niche interaction was higher or lower in one group compared to another.
@@ 2.6.3 Feature importance analysis
To interpret the MLP classifier, permutation importance (PI) [41, 42] was computed to quantify the contribution of each feature to model performance. For each feature, its values were randomly permuted across samples while keeping other features constant. The resulting decrease in weighted F1-score was recorded. This procedure was repeated multiple times to estimate the mean and standard deviation (SD) of the F1-score change.
@@ 3.2.2 Niche–gene features with information-theoretic selection
A total of 24 niche-gene features were identified (Table 2). HC-associated features were related to the genes PIGR, IGHG2, HLA-DRB1, IGHG1, CD38, STAT1, and CHI3L1 expressed in Niche 1, 2, and 4 cells; while UC-associated features were related to the genes IGKC, IGFBP5, HLA-DQB1, CASP3, and COL3A1 expressed in Niche 2, 3, and 4 cells; and CD-associated features were related to the genes MZT2A, COL9A2, FZD1, HDAC5, COL5A1, SOX6, FOXF1 expressed in Niche 3 and 4 cells. These features were integrated with niche composition and neighborhood enrichment features to form the final 44-feature dataset used for classification. Table 2 also indicates whether each gene has been previously reported in the literature to be associated with IBD. When marked “yes,” the corresponding reference cites the specific study supporting that association.
@@
Niche-Gene Feature MI Score Condition Group Adj. p-value Known IBD Association
niche_1_gene_HLA-DRB1 0.351 HC 0.037 yes [43]
niche_1_gene_IGHG2 0.351 HC <0.001 yes [44]
niche_1_gene_PIGR 0.357 HC <0.001 yes [45]
niche_2_gene_IGFBP5 0.271 UC <0.001 yes [46]
niche_2_gene_IGHG1 0.348 HC <0.001 yes [44]
niche_2_gene_IGHG2 0.352 HC <0.001 yes [44]
niche_3_gene_COL9A2 0.230 CD 0.020 no
niche_3_gene_HLA-DQB1 0.261 UC <0.001 yes [47]
niche_3_gene_IGFBP5 0.276 UC <0.001 yes [46]
niche_3_gene_MZT2A 0.244 CD <0.001 no
niche_4_gene_CASP3 0.280 UC <0.001 yes [48]
niche_4_gene_CD38 0.401 HC <0.001 yes [49]
niche_4_gene_CHI3L1 0.363 HC <0.001 yes [50]
niche_4_gene_COL3A1 0.267 UC <0.001 yes [51]
niche_4_gene_COL5A1 0.252 CD <0.001 yes [52]
niche_4_gene_FOXF1 0.211 CD 0.011 no
niche_4_gene_FZD1 0.352 CD <0.001 yes [53]
niche_4_gene_HDAC5 0.262 CD 0.004 yes [54]
niche_4_gene_HLA-DQB1 0.277 UC <0.001 yes [47]
niche_4_gene_IGHG2 0.369 HC <0.001 yes [44]
niche_4_gene_IGKC 0.313 UC 0.001 yes [55]
niche_4_gene_MZT2A 0.209 CD 0.002 no
niche_4_gene_SOX6 0.235 CD 0.010 yes [56]
niche_4_gene_STAT1 0.367 HC <0.001 yes [57]
@@ 3.3 Multilayer perceptron
Figure 5 shows the confusion matrices for both classification tasks, obtained by aggregating predictions across all three folds of group-stratified cross-validation (subjects were held out by group in each fold). In the three-class case (Figure 5(a)), HC samples were classified perfectly, while misclassifications primarily occurred between UC and CD, indicating overlapping features between these two disease states. In the two-class case (Figure 5(b)), the model showed strong separation between HC and IBD, with only a small number of misclassifications.
The corresponding classification reports, computed from the aggregated cross-validation results, are provided in Tables 3 and 4. Table 5 summarizes the mean and standard deviation of each performance metric across the three cross-validation folds, providing a measure of variability in model performance. For the more challenging three-class problem, the macro-averaged F1-score was 0.741 (accuracy $0.774\pm 0.161$), with perfect HC classification and errors between UC and CD, reflecting their biological similarity. In comparison, the two-class problem showed excellent separation of HC and IBD, with overall accuracy of $0.916\pm 0.118$.
Overall, these results indicate that while distinguishing between UC and CD remains challenging due to biological similarity, the MLP classifier is highly effective at differentiating IBD subjects from HC.
@@ 3.4.1 Causal discovery
Figure 6: Causal graphs depicting relationships between features and condition (Disease/Health State): (a) Niche composition, (b) Enrichment score, and (c) Niche-gene features. An open circle in an edge represents the potential presence of a latent confounder.
@@ 3.4.2 Statistical tests
Table 6: Pairwise comparisons of niche compositions across condition groups.
NMF Factor Group 1 Group 2 Difference Between Means Direction* p-value Observation
@@
Table 7: Pairwise comparisons of niche interactions across condition groups.